OpenStudy (konradzuse):

solve for y' = tan^3(3x)sec(3x)

OpenStudy (konradzuse):

y' = dy/dx right? so \[dy = \tan^3(3x)\sec(3x)dx\]

OpenStudy (anonymous):

tan^3(3x)sec(3x)=tan^2(3x)tan(3x)sec(3x)=(1+sec^2(3x)))tan(3x)sec(3x) put u=sec(3x) du = 3 sec(3x) tan(3x)

OpenStudy (anonymous):

\[ y=\int ( 1 +u^2) \frac {du}3 \]

OpenStudy (konradzuse):

hmm

OpenStudy (konradzuse):

ic hmm

OpenStudy (konradzuse):

\[1/3 [ u + u^3/3] +c\]

OpenStudy (konradzuse):

\[ u/3 + u^3/9 +c\]

OpenStudy (konradzuse):

sec(3x)/3 + sec^3(3x)/9 +c

OpenStudy (konradzuse):

I tihnk thats correct :D